[Anchal, Paco, JC]
We had to open ITMY, ETMY chamber doors to get the cavity aligned again. Once we did that, we regained cavity flashing and were able to align the input injection and cavity alignment to get transmission flashing to 1.0 (C1:LSC-TRY_OUT_DQ). JC later centered both ITMY and ETMY oplevs. The ITMY oplev had become completely out of range.
We also opened ITMX, ETMX chamber doors to get Xarm alignment. Again, it seems that ITMX had moved a lot due to cable post installation.
To be continued